OpenStudy (anonymous):

Find the General Solution with the given Initial Condition: y'+(2/t)y=(cost t)/t^2 y(pi)=0 t>0

OpenStudy (anonymous):

Integration Factor: \[\mu(t)=\int\limits\limits e ^{ \frac{ 2 }{ t }}\]

OpenStudy (anonymous):

i think you wrote it little bit different should be \[\Large I.F=e^{\int\limits_{}^{} \frac{2}{t}dt}\]

OpenStudy (anonymous):

ya that, but i just discovered that wolfram does these :)

OpenStudy (anonymous):

yes it does!

OpenStudy (anonymous):

medal for a medal lol:)

OpenStudy (anonymous):

it is false